Suppose that has a discrete uniform distribution on the integers 0 through 9. Determine the mean, variance, and standard deviation of the random variable . The mean is Enter your answer; The mean is [Round your answer to one decimal place (e.g. 98.7).] The variance is Enter your answer; The variance is [Round your answer to two decimal places (e.g. 98.76).] The standard deviation is Enter your answer; The standard deviation is [Round your answer to two decimal places (e.g. 98.76).]
Answer #1

Solution :

Given that,

a = 0

b = 9

mean = (a + b) / 2=(0+9)/2=4.5

mean=4.5

variance =  (b - a)2 / 12

= (9 - 0)2 / 12

=6.75

variance=6.75

standard deviation =  (b - a)2 / 12 = (9 - 0)2 / 12 = 2.60